OpenStudy (anonymous):

A train rolls past a stationary observer. To him, the train is moving at a speed of 25 m/s west, and a woman on the train is moving at a speed of 24.6 m/s west. How long does it take the woman to move 4 m relative to the train? A. 10 s B. 8.4 s C. 7.1 s D. 9.0 s

OpenStudy (anonymous):

The woman appears to move slower than the train. So she must be walking towards the back of the train with a velocity relative to the train of 25-24.6=0.4m/s So, to walk 4m relative to the train: time = distance/speed = 4/0.4 = 10s

OpenStudy (anonymous):

Thanks